if side of square field is 89 m.find the length of diagonal correct to 2 decimal places
Answer by nerdybill(7384)

if side of square field is 89 m.find the length of diagonal correct to 2 decimal places
.
apply Pythagorean theorem:
Let d = diagonal
then
d^2 = 89^2 + 89^2
d^2 = 7921 + 7921
d^2 = 15842
d = 125.87 m